OpenStudy (anonymous):

sin^(2) x = (1 - cos x)/2 ?

OpenStudy (anonymous):

\(\large sin^2x=\frac{1-cosx}{2} \) \(\large 1-cos^2x=\frac{1}{2}-\frac{cosx}{2} \) \(\large 0=cos^2x-\frac{cosx}{2}-\frac{1}{2} \) \(\large 0=2cos^2x-cosx-1 \) now it's a trig equation in quadratic form... can you solve from here?

OpenStudy (anonymous):

uhm..i just wanna know if sin ^(2) x is equal to (1-cosx)/2 according to the double angle formula??

OpenStudy (anonymous):

oh... you wanna know it the equation is an identity?

OpenStudy (anonymous):

i'll just google it myself...thanks for replying! :)